# Learning Algorithms Made Simple

### v1

@@ Abstract
In this paper, we discuss learning algorithms and their importance in different types of applications which includes training to identify important patterns and features in a straightforward, easy-to-understand manner. We will review the main concepts of artificial intelligence (AI), machine learning (ML), deep learning (DL), and hybrid models. Some important subsets of Machine Learning algorithms such as supervised, unsupervised, and reinforcement learning are also discussed in this paper. These techniques can be used for some important tasks like prediction, classification, and segmentation. Convolutional Neural Networks (CNNs) are used for image and video processing and many more applications. We dive into the architecture of CNNs and how to integrate CNNs with ML algorithms to build hybrid models. This paper explores the vulnerability of learning algorithms to noise, leading to misclassification. We further discuss the integration of learning algorithms with Large Language Models (LLM) to generate coherent responses applicable to many domains such as healthcare, marketing, and finance by learning important patterns from large volumes of data. Furthermore, we discuss the next generation of learning algorithms and how we may have an unified Adaptive and Dynamic Network to perform important tasks. Overall, this article provides brief overview of learning algorithms, exploring their current state, applications and future direction.
Index Terms:
Learning Algorithms, training, artificial intelligence (AI), classification, convolutional neural networks (CNNs), large language models (LLMs), patterns.
I Introduction
The definition of artificial intelligence is to use machines based on human intelligence to perform a task, such as decision-making or object recognition (see Fig. 1). Nowadays, AI technology and its applications are becoming very common in society, and are penetrating people’s lives in many ways. Its fast development is due to the rapid progress in science and technology [1]. We can claim that AI and its multidisciplinary applications will be categorized among the hottest subjects in the future. AI technologies can be seen everywhere, and in every technology-based institution such as Google, YouTube, Amazon, etc. It has vast applications in all fields of interest such as disease diagnosis, early prediction of tumors, recommended systems, image and signal processing, computer vision, etc. Experts can solve their problems more efficiently by employing AI compared to applying traditional methods and approaches. This technology is based on extensive computer programming for training machines with human beings’ behaviors for performing some important cognition-based tasks [2]. This is a full integration between humans and computer-based technologies [1]. Prediction and classification are categorized among the important tasks of learning-based systems. To better predict and classify, we are in need of suitable algorithms mimicking human behaviors, skills, and abilities to carry out and process these tasks [3]. As can be seen from Fig. 1, Machine Learning (ML) is a sub-field of AI, and Deep Learning (DL) is a sub-field of ML which we explained briefly in the following.
Figure 1: AI system and the relationship between AI, ML and DL[4].
Machine learning (ML) is a branch of artificial intelligence that creates and studies statistical algorithms that can learn from data and generalize to previously unknown data, allowing them to execute tasks without explicit instructions [5]. Training data is the sample data that a machine learning algorithm uses to build a mathematical model. This allows the system to execute future actions based on past information. Machine learning is widely utilized in many domains, such as computer vision, natural language processing, and robotics. It is essential in data mining, analysis, classification, and prediction. This technology is making our lives easier; for example, when we give a voice command on the phone or search for an image on the internet, it provides the results we need. Machine learning has numerous applications, including speech recognition, image recognition, and prediction. In speech recognition, software applications listen to human speech and transform it into text, simplifying many everyday tasks. Furthermore, the development of different algorithms has improved image recognition efficiency, allowing the system to comprehend an unknown image and deliver comprehensive information about it. In addition, the healthcare industry has emerged as a data-driven solution, enabling medical professionals to detect and categorize complicated diseases or find important patterns associated with a certain medical condition that would not be achievable through manual analysis.
Deep learning is a subfield of machine learning where neural networks are employed for performing various tasks. The origin of artificial neurons mimics the neurons of our nervous system, where a collection of many neurons is a neural network. This technique takes the dataset as input, passes the inputs through the neural network, extracts all the important features from the data, trains itself, and produces the output based on that. We can give images, text, and even sound as input. However, using this method requires a large amount of data to get good accuracy, because the system is trained with the training data and then tested with the testing data to see how effectively the system has learned. The time required for training using this approach depends on various parameters, such as the size of the network, the speed of the coding optimization, the processing unit, etc.
Deep learning has many important and successful applications that have given new dimensions to artificial intelligence. Researchers constantly explore deep learning as it recognizes complex patterns in images, text, words, and other data to generate accurate insights and predictions. Some of its notable applications are natural language processing, self-driving cars, object detection, virtual assistants like Google Assistant, and so on.
The future of learning algorithms seems promising, as researchers continue to work on and improve existing methods. However, Various issues must be addressed in the future to ensure the accuracy and strong performance of this technology. For example, because the healthcare industry is critical and patient-sensitive, developing a model for this field necessitates the availability of a high-quality dataset that can be utilized to train learning algorithms. Second, the interpretability and complexity of such learning algorithms raise several concerns when building a solution for a specific sector. Since different models have varying levels of complexity, they must be tuned to produce a fair and robust outcome. Another problem could be computer resources; for example, a deep learning model requires a large quantity of training data, and addressing these challenges may necessitate more lightweight algorithms and hardware developments.
@@ I Introduction
The rest of the paper is organized as follows. In Section II, we discuss various machine learning algorithms, including supervised learning, unsupervised learning, reinforcement learning, semi-supervised learning, federated learning, feature engineering/learning, transfer learning, and ensemble learning. In Section III, we explain deep learning algorithms such as CNNs, RNNs, LSTMs, GANs, the transformer architecture, and backpropagation. In Section IV, we demonstrate hybrid models. In Section V, we provide an explanation of the explainable artificial intelligence methods. In Section VI, we discuss fooling the learning algorithms into misclassification. In Section VII, we show how LLMs are integrated into various sectors. In Section VIII, we depicted real-world applications, and in Section IX, we explore the future of learning algorithms. Finally, in Section X, we wrap up the manuscript.
II Machine Learning
@@
Machine learning is categorized as a subset of artificial intelligence which deals with learning from the data to act like human beings, and to analyze and process the given tasks efficiently. It has a wide variety of applications in biomedical, engineering, science, agriculture, animal sciences, neuroscience, remote sensing, etc. Those equipped with ML can run decision-making and performance analysis much faster than they ever could have before. Segmentation, detection, and recognition are among the most important real-world problems. Segmentation is one of the critical steps to be done prior to further image analysis and processing. The segmented features are supposed to be fed to the algorithm for recognition and classification. Some published research has shown that ML can process these tasks properly [6]. Fig. 2 shows some of these tasks for Barbara test image [7]. In this section, we discuss some important concepts and techniques in ML.
@@ II-C Reinforcement Learning
II-D Semi-supervised Learning
@@
Figure 7: Illustrative frameworks of different machine learning paradigms: (a) supervised learning [16], which uses labeled data for training; (b) unsupervised learning [16], which aims to find patterns from unlabeled data; and (c) reinforcement learning [6], where an agent interacts with an environment to maximize cumulative rewards.
@@ II-F Feature Learning
We illustrated the feature learning paradigm in machine learning in Fig. 11. Under the paradigm shown in Fig.11, an initial (often sparse) feature set or raw data are input to learn implicit feature representations using a variety of techniques [26]. As a result, the feature representation is richer and frequently lower dimensional, which might improve performance when applied as the input to more specialized learning tasks [26].
@@ III Deep Learning
III-A Convolutional Neural Network (CNN)
@@
III-B Recurrent Neural Network (RNN)
@@
Recurrent networks are also distinguished by their parameter sharing across all network layers [34]. Unlike feedforward networks, which have distinct weights for each node, RNNs use the same weight parameter throughout all layers. Nevertheless, these weights are still modified using gradient descent and backpropagation techniques to support reinforcement learning [34].
To find the gradients (or derivatives), RNNs employ forward propagation and backpropagation through time (BPTT) methods. This approach differs slightly from standard backpropagation because it is tailored to sequence data [34]. The fundamentals of BPTT are similar to those of conventional backpropagation, where the model learns by estimating errors from its input layer to its output layer [34]. These computations enable us to suitably adjust the model’s parameters. The key difference with BPTT is that, while feedforward networks do not share parameters across layers, backpropagation through time (BPTT ) must accumulate errors at each time step [34].
III-C Long Short-Term Memory (LSTM)
@@
III-D Generative Adversarial Networks (GANs)
@@ III-E Transformer Architecture
III-F Backpropagation
The backpropagation algorithm calculates the gradient of a neural network’s loss function in terms of weights and biases [55]. It uses the chain rule to compute the partial derivatives of the loss function with respect to each weight and bias, starting from the output layer and working backward to the input layer (see Fig. 12.)
@@ III-F Backpropagation
$\frac{\delta C_{k}}{\delta W^{L}}=a^{L-1}\cdot\sigma^{\prime}(z^{L})\cdot 2(a^{L}-y)$ (7)
Here, $\frac{\delta C_{k}}{\delta W^{L}}$ denotes the partial derivative of the loss function $C_{k}$ with regard to weight $W^{L}$, assessing how the weight reduction responds to variations in that weight. This process iterates for each weight, bias, and activation number in the system. The average derivative [55] across data points is computed as:
$\frac{\delta C}{\delta W^{L}}=\frac{1}{n}\sum_{k=0}^{n-1}\frac{\delta C_{k}}{\delta W^{L}}$ (8)
@@
IV Hybrid Models
Although CNN itself is a powerful classification model but some believe that using hybrid CNN-ML models provides us with better results. As it is obvious, ConvNets can provide high dimensional vectors, and ML models like SVM can handle these complex decision boundaries well. So, in the hybrid-type models, CNN can be used to extract the meaningful features and during the training, these features are learned for some specific tasks. The features represent the input data (e.g. images). Each image is now a vector of learned features, and these features are positioned in the last convolutional layer before the fully connected layers. Here is an interesting question: will these learned features (which are in the last convolutional layer before fully connected layer) be fed directly to SVM? The layer that will be fed to SVM could either be the flatten layer which is the 1D representation of the features maps or the FC layer. 1D flatten layer represents the learned features obtained from last conv layer. If there will be a max pooling layer right before the flatten layer, then the number of neurons in flatten layer would be equal to the product of the dimensions of this max pooling layer. Thus, this flatten layer will be fed into SVM as an input. The extracted features are fed into ML model like SVM for further classification. SVM is trained based on these learned features extracted from CNN. SVM uses these features to learn how to perform classification. Since it is a hybrid model, there may be a question of whether or not these extracted features are already learned, or they will be learned once they are fed into SVM? So, the extracted features from CNN are already learned during training. Once they are fed into SVM, then it uses these features to learn how to perform the classification and prediction. Note that the SVM is not trained already. It will be trained using these extracted features. So there will be two training phases in hybrid models. One is CNN training (based on the input data) and the other is SVM training (based on the extracted features from CNN).
V Explainable Artificial Intelligence (XAI)
@@
VI Fooling Learning Algorithms into Misclassification
An adversarial attack is a term used to describe a system designed to fool learning-based models into making mistakes (misclassifying an image). They exploit weaknesses in the models and take advantage of properties such as shortcutting in order to subtly alter an image just enough to make the model classify an image differently. Modern machine learning models are susceptible to these attacks, as they have inherent limitations and vulnerabilities, ranging from the data they are trained on to their parameters and lack of transparency. Through these multiple limitations, adversarial attacks can be created with surprising ease, and go on to become very effective and discreet, as they are often not detectable by simply looking at the images. Attack methods such as PGD and FGSM are examples of this. Even high-performing state-of-the-art models can be very susceptible to attack. It is extremely important to study adversarial noise removal and work to eliminate malicious attempts by developing robust noise-informed learning algorithms, as successful attacks could lead to disastrous consequences in many fields, such as healthcare or autonomous driving. For example, as can be seen from Fig. 13, adversarial noise affected the cancer edges by adding more pixels around them, which may lead to misclassifying the cancerous image into non-cancerous. The dataset is available in [57].
Figure 13: Effect of adversarial noise on the image. (a) clean image, (b) clean edge, (c) noisy image, (d) noisy edge.
@@ VII LLM Integration
Table I provides an overview of LLM applications across healthcare, education, security, language processing, business, and content creation. In the healthcare sector, these models have been adopted to automate administrative tasks and develop Question answering (QA) systems. Hospitals, which traditionally require significant manpower for administrative duties, are now automating these processes using LLMs like GPT and Llama. Additionally, Question-answering (QA) frameworks allow patients to ask questions and receive automated responses without waiting for a human agent.
@@
Looking at Table I, we can also see that researchers have been very interested in fine-tuning LLMs; very few studies have not addressed this topic. Furthermore, deployment was not the main focus of the previous research listed in Table I, which meant that the models were maintained in the testing stage, where they are still undergoing constant improvement. In addition, a group of researchers [59] concentrates on both deployment and fine-tuning because an end-to-end solution was considered while implementing a tailored chatbot to access university resources.

### v3

# AI Learning Algorithms: Deep Learning, Hybrid Models, and Large-Scale Model Integration

@@ Abstract
In this paper, we discuss learning algorithms and their importance across various applications, emphasizing their ability to identify key patterns and features in a straightforward, easy-to-understand manner. We review the main concepts of Artificial Intelligence (AI), Machine Learning (ML), Deep Learning (DL), and hybrid models. Key subsets of ML algorithms, such as supervised, unsupervised, and reinforcement learning, are also covered. These techniques are essential for tasks such as prediction, classification, and segmentation. Convolutional Neural Networks (CNNs), widely used in image and video processing, are explored in detail, including their architecture and how they can be integrated with ML algorithms to build hybrid models. This paper also explores the vulnerability of learning algorithms to adversarial noise, which can lead to misclassification. Furthermore, we discuss the integration of learning algorithms with Large Language Models (LLMs) to generate coherent responses in domains such as healthcare, marketing, and finance by learning and extracting meaningful patterns from large datasets. Finally, we discuss the next generation of learning algorithms, emphasizing the potential for a unified, adaptive, and dynamic network capable of performing a wide range of tasks. Overall, this article provides a concise overview of learning algorithms, highlighting their current capabilities, applications, and future directions.
{IEEEkeywords}
Learning algorithms, AI, deep learning, hybrid models, LLM, patterns.
1 Introduction
Learning algorithms are computational techniques that autonomously refine their performance by accumulating experience, detecting patterns within data, and making informed predictions or decisions based on that data. These algorithms are integral components of Artificial Intelligence (AI), Machine Learning (ML) and Deep Learning (DL) frameworks.
Artificial Intelligence (AI) refers to the use of machines to perform tasks traditionally requiring human intelligence, such as decision-making or object recognition (see Fig. 1). In recent years, AI technologies have rapidly evolved, becoming deeply embedded in everyday life. This growth has been fueled by significant advancements in science and technology [1], positioning AI and its multidisciplinary applications among the most influential subjects of the future. AI is now ubiquitous across major technology-driven platforms such as Google, YouTube, and Amazon. It has found widespread use in domains including disease diagnosis, early tumor prediction, recommendation systems, image and signal processing, and computer vision. Unlike traditional approaches, AI empowers experts to solve complex problems more efficiently by leveraging extensive computer programming that enables machines to emulate human behavior [2]. This represents a full integration between human cognition and computer-based technologies [1].
Prediction and classification are among the most essential cognitive tasks that AI-based systems aim to replicate. Achieving high accuracy in these tasks requires sophisticated algorithms capable of mimicking human skills and decision-making abilities [3]. As illustrated in Fig. 1 [5], Machine Learning (ML) is a sub-field of AI, while Deep Learning (DL) is a further subset within ML. These learning paradigms form the computational backbone of modern AI systems. Brain-inspired AI approaches, particularly Deep Learning and Cognitive Neuroscience, further refine AI’s ability to process information similarly to the human brain (see Fig. 1). These approaches leverage neural network architectures that emulate biological neurons, enhancing AI’s ability to tackle complex, real-world problems efficiently. By adopting strategies inspired by the brain’s structure and functionality, these methods enable AI systems to exhibit adaptive learning, robust pattern recognition, and enhanced problem-solving capabilities, making them a crucial advancement in intelligent system design.
Figure 1: AI system and the relationship between AI, ML and DL.
ML involves developing statistical models that can learn from data and generalize to unseen data, enabling systems to execute tasks without explicit instructions [6]. It plays a crucial role in fields like computer vision, natural language processing, and robotics. The concept of training data—data used to build a predictive model—is foundational to ML systems. ML applications span from speech and image recognition to personalized recommendation systems and predictive analytics in healthcare.
DL, a specialized branch of ML, employs neural networks to perform tasks by automatically learning hierarchical feature representations from large datasets. Inspired by the structure of biological neurons, DL models process inputs such as text, images, or audio through deep architectures. The accuracy and effectiveness of DL models are influenced by factors such as dataset size, network architecture, optimization techniques, and computational resources. DL has been instrumental in advancing technologies like natural language processing, autonomous driving, object detection, and virtual assistants.
Despite their success, learning algorithms still face challenges. In critical domains like healthcare, high-quality annotated datasets are essential. Moreover, the interpretability and complexity of these models can hinder their adoption. As DL models often require substantial computational resources, future developments may focus on lightweight algorithms and efficient hardware solutions to improve scalability and accessibility.
Although numerous review articles have explored learning algorithms [116] [117] [118], most focus exclusively on either classical machine learning techniques or deep learning models. In contrast, this review provides an integrative and up-to-date analysis—ranging from traditional ML and DL to emerging paradigms such as Explainable AI, adversarial robustness, federated learning, and Large Language Models. This comprehensive scope enables readers to understand both foundational concepts and cutting-edge trends within a single framework, making this review relevant to both academic and applied research communities.
This article aims to provide a systematic and unified overview of the wide range of learning algorithms available in the AI domain. As technology advances rapidly and domain-specific expertise becomes more widespread, researchers often face fragmented knowledge. This review integrates various learning paradigms to serve as a comprehensive resource for both early-stage learners and experienced researchers. It also illustrates how learning algorithms change and adapt in practical applications by discussing both established and new directions.
The key contributions of this article have been outlined as follows:
  • A comprehensive overview of both classical and modern learning algorithms is provided in a straightforward manner, with examples that help unify the scattered knowledge in the field.
  • A systematic comparison between traditional machine learning paradigms and contemporary approaches, including deep learning and hybrid models, has been conducted.
  • Emerging topics such as Explainable AI (XAI), adversarial robustness, federated learning, and Large Language Models (LLMs) have been incorporated to reflect current advancements.
  • Fragmented knowledge across different domains has been bridged by integrating various learning paradigms into a cohesive review.
  • Practical applications in sensitive domains, such as healthcare and natural language understanding, have been emphasized to highlight real-world impact.
@@ 1 Introduction
The rest of the paper is organized as follows. In Section 2, we discuss various machine learning algorithms, including supervised learning, unsupervised learning, reinforcement learning, semi-supervised learning, federated learning, feature engineering/learning, transfer learning, and ensemble learning. In Section 3, we explain deep learning algorithms such as Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Long Short-Term Memory networks (LSTMs), Generative Adversarial Networks (GANs), the transformer architecture, and backpropagation. In Section 4, we demonstrate hybrid models. In Section 5, we provide an explanation of the explainable artificial intelligence methods. In Section 6, we discuss fooling the learning algorithms into misclassification. In Section 7, we show how LLMs are integrated into various sectors. In Section 8, we depict real-world applications, and in Section 9, we explore the future of learning algorithms. Finally, in Section 10, we wrap up the manuscript.
2 Machine Learning
@@
Machine learning is categorized as a subset of artificial intelligence that encompasses probability theory, statistics, approximation theory, convex analysis, and algorithm complexity which deals with learning from the data to act like human beings, and to analyze and process the given tasks efficiently. It has a wide variety of applications in biomedical, engineering, science, agriculture, animal sciences, neuroscience, remote sensing, etc. Those equipped with ML can run decision-making and performance analysis much faster than they ever could have before. However, this section is dedicated to some fundamental computer vision tasks like segmentation, classification, and detection, as they are highly applicable in numerous practical scenarios.
Segmentation is one of the critical steps to be done prior to further image analysis and processing. The segmented features are supposed to be fed to the algorithm for recognition and classification. Some published research has shown that ML can process these tasks properly [7]. Fig. 2 shows some of these tasks for Barbara test image [8]. In this section, we also discuss some important concepts and techniques in ML.
@@ 2.3 Reinforcement Learning
2.4 Semi-supervised Learning
@@
Figure 7: Illustrative frameworks of different machine learning paradigms: (a) supervised learning, which uses labeled data for training; (b) unsupervised learning, which aims to find patterns from unlabeled data; and (c) reinforcement learning [7], where an agent interacts with an environment to maximize cumulative rewards.
@@ 2.6 Feature Learning
We illustrated the feature learning paradigm in machine learning in Fig. 11. As it is shown in this figure, an initial (often sparse) feature set or raw data are input to learn implicit feature representations using a variety of techniques [27]. As a result, the feature representation is richer and frequently lower dimensional, which might improve performance when applied as the input to more specialized learning tasks [27].
@@ 3 Deep Learning
3.1 Convolutional Neural Network (CNN)
@@
3.2 Recurrent Neural Network (RNN)
@@
Recurrent networks are also distinguished by their parameter sharing across all network layers [35]. Unlike feedforward networks, which have distinct weights for each node, RNNs use the same weight parameter throughout all layers. Nevertheless, these weights are still modified using gradient descent and backpropagation techniques to support reinforcement learning [35]. To find the gradients (or derivatives), RNNs employ forward propagation and backpropagation through time (BPTT) methods. This approach differs slightly from standard backpropagation because it is tailored to sequence data [35]. The fundamentals of BPTT are similar to those of conventional backpropagation, where the model learns by estimating errors from its input layer to its output layer [35]. These computations enable us to suitably adjust the model’s parameters. The key difference with BPTT is that, while feedforward networks do not share parameters across layers, backpropagation through time (BPTT ) must accumulate errors at each time step [35].
3.3 Long Short-Term Memory (LSTM)
@@
3.4 Generative Adversarial Networks (GANs)
@@ 3.5 Transformer Architecture
3.6 Backpropagation
The backpropagation algorithm calculates the gradient of a neural network’s loss function in terms of weights and biases [56]. It uses the chain rule to compute the partial derivatives of the loss function with respect to each weight and bias, starting from the output layer and working backward to the input layer (see Fig. 12).
@@ 3.6 Backpropagation
$\frac{\partial C_{k}}{\partial W^{L}}=a^{L-1}\cdot\sigma^{\prime}(z^{L})\cdot 2(a^{L}-y)$ (7)
Here, $\frac{\partial C_{k}}{\partial W^{L}}$ denotes the partial derivative of the loss function $C_{k}$ with regard to weight $W^{L}$, assessing how the weight reduction responds to variations in that weight. This process iterates for each weight, bias, and activation number in the system. The average derivative [56] across data points is computed as:
$\frac{\partial C}{\partial W^{L}}=\frac{1}{n}\sum_{k=0}^{n-1}\frac{\partial C_{k}}{\partial W^{L}}$ (8)
@@
4 Hybrid Models
Although CNN itself is a powerful classification model but some believe that using hybrid CNN-ML models provides us with better results. As it is obvious, ConvNets can provide high dimensional vectors, and ML models like SVM can handle these complex decision boundaries well. So, in the hybrid-type models, CNN can be used to extract the meaningful features and during the training, these features are learned for some specific tasks (see Fig. 16). The features represent the input data (e.g. images). Each image is now a vector of learned features, and these features are positioned in the last convolutional layer before the fully connected layers. Here is an interesting question: will the features learned in the last convolutional layer—just before the fully connected layer—be fed directly into the SVM? The layer that will be fed to SVM could either be the flatten layer which is the 1D representation of the features maps or the FC layer. 1D flatten layer represents the learned features obtained from last conv layer. If there will be a max pooling layer right before the flatten layer, then the number of neurons in flatten layer would be equal to the product of the dimensions of this max pooling layer. Thus, this flatten layer will be fed into a ML model like SVM as an input for further classification. The SVM is trained using features extracted from the CNN. These features, learned during CNN training, serve as input to the SVM, which then learns how to perform classification. Since this is a hybrid model, an important consideration is whether the features are learned before or during SVM training. In fact, the features are already learned during the CNN training phase. The SVM is trained separately using these extracted features. Therefore, hybrid models like this involve two training stages: first, the CNN is trained on the input data to extract features, and then the SVM is trained on those features to perform classification and prediction.
5 Explainable Artificial Intelligence (XAI)
@@
6 Fooling Learning Algorithms into Misclassification
An adversarial attack is a term used to describe a system designed to fool learning-based models into making mistakes (misclassifying an image). They exploit weaknesses in the models and take advantage of properties such as shortcutting in order to subtly alter an image just enough to make the model classify an image differently. Modern machine learning models are susceptible to these attacks, as they have inherent limitations and vulnerabilities, ranging from the data they are trained on to their parameters and lack of transparency. Through these multiple limitations, adversarial attacks can be created with surprising ease, and go on to become very effective and discreet, as they are often not detectable by simply looking at the images. Examples include attack methods such as Projected Gradient Descent (PGD) and Fast Gradient Sign Method (FGSM). Even high-performing state-of-the-art models can be very susceptible to attack. It is extremely important to study adversarial noise removal and work to eliminate malicious attempts by developing robust noise-informed learning algorithms, as successful attacks could lead to disastrous consequences in many fields, such as healthcare or autonomous driving. For example, as can be seen from Fig. 13, adversarial noise affected the cancer edges by adding more pixels around them, which may lead to misclassifying the cancerous image into non-cancerous. The dataset is available in [58].
Figure 13: Effect of FGSM adversarial noise on the image. (a) clean image, (b) clean edge, (c) noisy image, (d) noisy edge.
@@ 7.1 How LLMs Power Real-World Applications?
Table 1 provides an overview of LLM applications across healthcare, education, security, language processing, business, and content creation. In the healthcare sector, these models have been adopted to automate administrative tasks and develop Question answering (QA) systems. Hospitals, which traditionally require significant manpower for administrative duties, are now automating these processes using LLMs like GPT and Llama. Additionally, Question-answering (QA) frameworks allow patients to ask questions and receive automated responses without waiting for a human agent.
@@
Looking at Table 1, we can also see that researchers have been very interested in fine-tuning LLMs; very few studies have not addressed this topic. Furthermore, deployment was not the main focus of the previous research listed in Table 1, which meant that the models were maintained in the testing stage, where they are still undergoing constant improvement. In addition, a group of researchers [60] concentrates on both deployment and fine-tuning because an end-to-end solution was considered while implementing a tailored chatbot to access university resources.
7.2 LLM Reasoning
LLM reasoning is the ability of large language models to imitate human-like cognitive processes when addressing problems requiring multi-step logic, deduction, or inference. LLMs that can solve mathematical problems, complete code, and conduct multi-step logical inference show remarkable reasoning abilities. However, models like as GPT-4 and LLaMA3 continue to struggle with accuracy in complicated, multi-step reasoning, where surface-level correctness does not ensure logical consistency.
To address this, recent work by Zhenwen et al. [111] proposes a verification framework to enhance inference-time reasoning through verifiers such as Math-Rev and Code-Rev. These are trained using SimPO—a reference-free preference optimization technique—to distinguish correct from incorrect outputs sampled from various LLMs. The authors also introduce CoTnPoT, a hybrid technique that transforms natural language reasoning steps $S_{\text{CoT}}$ into executable programs $S_{\text{PoT}}$:
$S_{\text{PoT}}=\text{CoderLLM}(Q,S_{\text{CoT}})$ (9)
where $Q$ is the original problem prompt, $S_{\text{CoT}}$ is the chain-of-thought solution in natural language, and $\text{CoderLLM}(\cdot)$ is a coder model that turns reasoning stages into executable code-like representations for validation.
During inference, solutions are scored using Gumbel-Softmax weighted voting as follows:
$y_{i}=\frac{\exp\left(\frac{\log(\pi_{i})}{\tau}\right)}{\sum_{j=1}^{k}\exp\left(\frac{\log(\pi_{j})}{\tau}\right)}$ (10)
where $\pi_{i}$ is the unnormalized score awarded by the verifier to the $i$-th solution, $\tau$ is the temperature parameter determining the sharpness of the distribution, and $y_{i}$ is the normalized weight utilized for final decision-making.
However, this approach improves reasoning accuracy on datasets like GSM8k [114] and MathBench [115], outperforming even GPT-4o when paired with strong backbone models. Building upon the need for standardized evaluation, Hao et al. [112] propose AutoRace, an automated reasoning chain evaluator that identifies common reasoning errors without relying on handcrafted prompts. It defines a binary metric $s(z)\in\{0,1\}$ based on whether an LLM-generated reasoning chain $z$ leads to a correct answer $y$ matching the reference $y_{r}$. Alongside this, the authors introduce LLM Reasoners, a modular library unifying reasoning algorithms via a search-based formulation, which is defined as follows:
$\arg\max_{a_{0},\ldots,a_{T}}\sum_{t=0}^{T}r(s_{t},a_{t}),\quad s_{t}\sim T(\cdot|s_{t-1},a_{t-1})$ (11)
where $a_{t}$ is the action performed at step $t$, $s_{t}$ is the reasoning state that results, $r(s_{t},a_{t})$ is the reward function that assesses the quality of each reasoning step, and $T(\cdot|s_{t-1},a_{t-1})$ is the world model that defines the state transitions based on earlier actions.
This framework supports consistent comparison of CoT, Tree-of-Thought (ToT), and Reasoning via Action Planning (RAP), with RAP showing superior performance by combining explicit world models with reward-guided search (e.g., Monte Carlo Tree Search). Complementary to these structural and evaluative advances, Han et al. [113] address the inefficiency of CoT-style reasoning, which often incurs excessive token usage. They introduce TALE—a Token-budget-Aware LLM reasoning framework—to reduce costs while preserving performance. TALE includes two methods: TALE-EP, which uses zero-shot token budget estimation to create concise prompts, and TALE-PT, which trains LLMs to internalize token-efficient reasoning via Supervised Fine-Tuning (SFT) or Direct Preference Optimization (DPO). These are trained using objectives and we can represent them mathematically as follows:
$\mathcal{L}_{\text{CE}}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}\log P(y_{i,t}\mid y_{i,<t},x_{i})$ (12)
where $\mathcal{L}_{\text{CE}}$ denotes the cross-entropy loss over the training set, $y_{i,t}$ is the target token at time step $t$, $y_{i,<t}$ is the sequence of previous tokens, and $x_{i}$ is the input question for the $i$-th example.
$\mathcal{L}_{\text{DPO}}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(s(y_{i},x_{i}))}{\exp(s(y_{i},x_{i}))+\exp(s(y^{\prime}_{i},x_{i}))}$ (13)
where $s(y,x)$ represents the log-likelihood score of a candidate output $y$ given input $x$, and $y_{i}\succ y^{\prime}_{i}$ indicates that the model should prefer the token-efficient output $y_{i}$ over the baseline output $y^{\prime}_{i}$.
This preference-based loss function encourages alignment with concise and accurate responses. Results on benchmarks: GSM8k [114] and MathBench [115] reveal that TALE achieves up to 68% reduction in token usage and 59% in cost, with minimal accuracy loss. Together, these efforts advance the field of LLM reasoning by tackling performance, evaluation, and efficiency, laying the groundwork for more accurate, interpretable, and cost-effective reasoning systems.